Production Report Tomar Company produces vitamin energy drinks. The Mixing Department, the first process department, mixes the ingredients required for the drinks. The following data are for April: Work in process, April 1 — Quarts started 90,000 Quarts transferred out 75,000 Quarts in EWIP 15,000 Direct materials cost $282,000 Direct labor cost $167,000 Overhead applied $334,000 Direct materials are added throughout the process. Ending inventory is 80 percent complete with respect to direct labor and overhead. Required: Prepare a production report for the Mixing Department for April. If an answer is zero, enter "0". Tomar Company Mixing Department Production Report for April Unit Information Units to account for: Total units to account for Units accounted for: Physical Flow Equivalent Units Total units accounted for Work completed Cost Information Costs to account for: $ Total costs to account for $ Cost per equivalent unit $ Costs accounted for: $ Total costs accounted for $
Answer #1
Tomar Company
Mixing Department
Production Report for April
Unit Information
Units to account for:
Units in beginning work in process 0
Units started 90000
Total units to account for 90000
Physical Flow Equivalent Units
Units accounted for:
Units completed 75000 75000
Units in ending work in process 15000 12000
Total Units accounted for 90000
Work completed 87000
Cost information
Costs to account for:
Beginning work in process 0
Incurred during the period 783000 =282000+167000+334000
Total Costs to account for 783000
Divided by Equivalent units 87000
Cost per equivalent unit 9.00
Costs accounted for:
Goods transferred out 675000 =75000*9
Ending work in process 108000 =12000*9
Total Costs accounted for 783000